OpenStudy (anonymous):

deteminate the polynomial if so it gets a factor x-k x^2-7x+12

OpenStudy (lgbasallote):

are you familiar with remainder theorem?

OpenStudy (lgbasallote):

i guess nobody's home

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

x - k is a factor therefore p(k) = 0 p(x) = x^2 - 7x + 12 p(k) = k^2 - 7k + 12 therefore we just equate that to 0 k^2 - 7k + 12 = 0 (k-4)(k-3) = 0 therefore your possible roots are k = 4 and k = 3

OpenStudy (lgbasallote):

does that help/

OpenStudy (anonymous):

thanks =)